find the sum of series 101+99+97+..+47
Answer by rothauserc(4718)

101+99+97+..+47
this is an arithmetic series, the nth term is found by
Xn = 101 -(2 * (n-1))
*****************************************************
we need to determine how many terms to sum
47 = 101 -(2*(n-1))
47 = 101 -2n +2
2n = 56
n = 28
there are 28 terms to sum
*****************************************************
sum = (28/2)*(101+47)
sum = 14 * 148
sum = 2072